# T-araVLN: Translator for Agricultural Robotic Agents on Vision-and-Language Navigation

### v2

@@ Abstract
Agricultural robotic agents have been becoming powerful helpers in a wide range of agricultural tasks, nevertheless, still heavily rely on manual operation or untransportable railway for movement. The AgriVLN method and the A2A benchmark pioneeringly extend Vision-and-Language Navigation (VLN) to the agricultural domain, enabling agents navigate to the target position following the natural language instructions. AgriVLN effectively understands the simple instructions, however, often misunderstands the complicated instructions. To bridge this gap, we propose the method of Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), in which the Instruction Translator module translates the original instruction to be both refined and precise. Being evaluated on the A2A benchmark, our T-araVLN effectively improves SR from 0.47 to 0.63 and reduces NE from 2.91m to 2.28m, demonstrating the state-of-the-art performance in the agricultural domain. Code: https://github.com/AlexTraveling/T-araVLN
@@ 1 Introduction
The most similar study to ours is VLN-Trans [14], which trains the translator module to convert original instructions into easy-to-follow sub-instruction representations, effectively solving two common types of instruction issue. However, the instruction issue types in A2A are much more complicated than in other mainstream VLN benchmarks [5, 15, 16], which makes training-based methods like VLN-Trans [14] and MTST [17] struggling with the circumstances out of training data, leaving a big issue on generalization capability.
@@ 3.2 Qualitative Experiment
# Method Translator A2A (subtask == 2) A2A (subtask $\geq$ 3) A2A
@@
Table 1: Comparison Experiment results of baseline models and our T-araVLN on the two-subtasks portion, the no-less-than-three-subtasks portion and the whole of A2A benchmark. Bold and underline represent the best and worst scores, respectively.
@@ 3.3 Comparison Experiment
The results of Comparison Experiment is shown in Table 1. On the complete version of A2A, our T-araVLN with GPT-4.1-based translator (#8) achieves SR of 0.63 and NE of 2.28, surpassing all the three baselines (#3, #4 and #5) on both SR and NE. On the two-subtasks portion, our T-araVLN even improves SR from 0.58 to 0.80, which nearly approaches the performance of human. In summary, Comparison Experiment proves the significant effectiveness of the Instruction Translator module, which leads to the state-of-the-art performance of T-araVLN in the agricultural VLN domain.
@@ 3.4.1 Large Language Model
We test T-araVLN with two more different LLMs, as shown in Table 1. When the Instruction Translator module is driven by Claude-3.7-sonnet (#6) or DeepSeek-r1 (#7), the BERTScores are 0.9105 and 0.8980, demonstrating more conservative and more radical translation strategies, respectively. However, SR decrease by 5 or 4 percentage points and NE increase by 0.69 or 0.27 meters, respectively, which proves that overly conservation may bring tedious noises and overly radical may delete essential semantics. Therefore, we select GPT-4.1 as the LLM for the Instruction Translator module in T-araVLN.
@@ 3.4.2 Translation Principle
We ablate the five principles in the module of Instruction Translator, in which we follow GPT-4.1 as LLM, as shown in Table 2 and Figure 4. When single principle is applied (#11, #12, #13, #14 or #15), the performance improves by 2 $\sim$ 10 percentage points on SR and 0.01 $\sim$ 0.52 meters on NE, which indicates the effectiveness of every principle. When all principles are combined (#16), the performance improves to the best, i.e., our proposed T-araVLN, which indicates that every principle contributes to different aspects of translation, meanwhile, all principles collaborate without significant interference.
@@
Table 2: Ablation Experiment results on Translation Principle in translator. “+” represents integration. Bold and underline represent the best and worst scores, respectively.
@@
Fig. 4: Ablation Experiment illustration on Translation Principle in translator. “+” represents integration.
@@ 4 Conclusion
In this paper, we propose Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), in which the Instruction Translator module translates the original instruction to be both refined and precise. Being evaluated on the A2A benchmark, our T-araVLN effectively improves SR from 0.47 to 0.63 and reduces NE from 2.91m to 2.28m, demonstrating the state-of-the-art performance in the agricultural domain.

### v3

@@ Abstract
Agricultural robotic agents have been becoming powerful helpers in a wide range of agricultural tasks, nevertheless, still heavily rely on manual operation or untransportable railway for movement. The AgriVLN method and the A2A benchmark pioneeringly extend Vision-and-Language Navigation (VLN) to the agricultural domain, enabling agents navigate to the target position following the natural language instructions. AgriVLN effectively understands the simple instructions, however, often misunderstands the complicated instructions. To bridge this gap, we propose the method of Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), in which the Instruction Translator module translates the original instruction to be both refined and precise. Being evaluated on the A2A benchmark, our T-araVLN effectively improves Success Rate from 0.47 to 0.63 and reduces Navigation Error from 2.91m to 2.28m, demonstrating the state-of-the-art performance in the agricultural domain. Code: https://github.com/AlexTraveling/T-araVLN
@@ 1 Introduction
The most relevant study to ours is VLN-Trans [14], which trains the translator module to convert original instructions into easy-to-follow sub-instruction representations, effectively solving two common types of instruction issue. However, the instruction issue types in A2A are much more complicated than in other mainstream VLN benchmarks [5, 15, 16], which makes training-based methods like VLN-Trans [14] and MTST [17] struggling with the circumstances out of training data, leaving a big issue on generalization capability.
@@ 3.2 Qualitative Experiment
# Method Translator A2A ( low-complexity ) A2A ( high-complexity ) A2A
@@
Table 1: Comparison Experiment results between T-araVLN and state-of-the-art methods on the low-complexity portion (subtask == 2), the high-complexity portion (subtask $\geq$ 3), and the whole of A2A. Bold and underline represent the best and worst scores, respectively.
@@ 3.3 Comparison Experiment
The results of Comparison Experiment is shown in Table 1. On the whole of A2A, our T-araVLN with GPT-4.1-based translator (#8) achieves SR of 0.63 and NE of 2.28, surpassing all the three baselines (#3, #4 and #5) on both SR and NE. On the low-complexity portion, our T-araVLN even substantially improves SR from 0.58 to 0.80, which nearly approaches the performance of human. In summary, Comparison Experiment proves the significant effectiveness of the Instruction Translator module, which leads to the state-of-the-art performance of T-araVLN in the agricultural VLN domain.
@@ 3.4.1 Large Language Model
We test T-araVLN with two additional LLMs of comparable size and performance, as shown in Table 1. When the Instruction Translator module is driven by Claude-3.7-sonnet (#6) or DeepSeek-r1 (#7), the BERTScores are 0.9105 and 0.8980, demonstrating more conservative and more radical translation strategies, respectively. However, SR decrease by 5 or 4 percentage points and NE increase by 0.69 or 0.27 meters, respectively, which suggests that overly conservation may bring tedious noises and overly radical may delete essential semantics. Compared to them, GPT-4.1 achieves the best trade-off. Therefore, we select GPT-4.1 as the LLM for the Instruction Translator module in T-araVLN.
@@ 3.4.2 Translation Principle
We ablate the five Translation Principles in the module of Instruction Translator, in which we follow GPT-4.1 as LLM, as shown in Table 2 and Figure 4. When every single principle is applied (#11, #12, #13, #14 or #15), the performance improves by 2 $\sim$ 10 percentage points on SR and 0.01 $\sim$ 0.52 meters on NE, which indicates the effectiveness of every principle. When all principles are combined (#16), the performance improves to the best, i.e., our proposed T-araVLN, which indicates that every principle contributes to a different emphasis of translation, meanwhile, all principles collaborate without significant interference.
@@
Table 2: Ablation Experiment results on the five Translation Principles in Instruction Translator. “+” represents integration. Bold and underline represent the best and worst scores, respectively.
@@
Fig. 4: Ablation Experiment illustration on the five Translation Principles in Instruction Translator. “+” represents integration.
@@ 4 Conclusion
In this paper, we propose Translator for Agricultural Robotic Agents on Vision-and-Language Navigation (T-araVLN), in which the Instruction Translator module translates the original instruction to be both refined and precise. Being evaluated on the A2A benchmark, our T-araVLN effectively improves Success Rate from 0.47 to 0.63 and reduces Navigation Error from 2.91m to 2.28m, demonstrating the state-of-the-art performance in the agricultural domain.